When conducting a t test for the correlation coefficient in a study with 50 individuals, the degrees of freedom will be

  1. 48  
  2. 49  
  3. 51  
  4. 52  
Answer #1

Sample size =n=50

Degree of freedom for hypothesis test of correlation =n-2=50-2=48

A) 48